### v1

# GAST: Sequential Gaussian Avatars with Hierarchical Spatio-temporal Context

Abstract
3D human avatars, through the use of canonical radiance fields and per-frame observed warping, enable high-fidelity rendering and animating. However, existing methods, which rely on either spatial SMPL(-X) poses or temporal embeddings, respectively suffer from coarse rendering quality or limited animation flexibility. To address these challenges, we propose GAST, a framework that unifies 3D human modeling with 3DGS by hierarchically integrating both spatial and temporal information. Specifically, we design a sequential conditioning framework for the non-rigid warping of the human body, under whose guidance more accurate 3D Gaussians can be obtained in the observation space. Moreover, the explicit properties of Gaussians allow us to embed richer sequential information, encompassing both the coarse sequence of human poses and finer per-vertex motion details. These sequence conditions are further sampled across different temporal scales, in a coarse-to-fine manner, ensuring unbiased inputs for non-rigid warping. Experimental results demonstrate that our method combined with hierarchical spatio-temporal modeling surpasses concurrent baselines, delivering both high-quality rendering and flexible animating capabilities.
1 Introduction
Figure 1: Illustration of the Hierarchical Spatio-temporal Context Design. We model the motion-dependent appearance variations with both coarse skeleton condition $\Delta P$ and fine-grained point-wise velocity condition $\mathcal{V}$. $\Delta P$ describes the overall human pose motion, which is derived from the difference between the human poses at adjacent frames. $\mathcal{V}$ is the point-wise velocity that indicates finer-grained motion in local regions. Based on $\Delta P$ and $\mathcal{V}$, we utilize an MLP to predict more accurate non-rigid deformation of each Gaussian primitive. To achieve robust deformation, we sample the overall motion trend and inter-frame details via diverse time intervals $s$ for $\Delta P$ and $\mathcal{V}$. The rendering results demonstrate that our method excels in regions with complex texture details as well as in areas where the appearance varies dramatically with motions.
Recent research on digital humans has highlighted the efficiency of 3D Gaussian Splatting [23] (3DGS), demonstrating its capability for high-quality and real-time rendering. By defining T-pose human Gaussians in a canonical space and employing specific warping techniques, we can render human avatars under various observations. The canonical Gaussian units and warping weights are jointly optimized under supervision from human video input.
The primary bottleneck in this task stems from the limited modeling of non-rigid warping, requiring a trade-off between rendering quality and animating capability. One stream of approaches [17, 44] introduces the prior from SMPL(-X) [37, 42] poses, leveraging human poses to guide motions of each observation and relying on Linear Blend Skinning (LBS) for motion warping. Although they are well-suited for pose-driven animation, they often struggle with per-frame non-rigid warping in scenarios with complex garments. Other approaches [33, 51] utilize temporal embeddings instead of SMPL poses to model each frame independently, achieving high-quality rendering but making it challenging to animate avatars.
We experimentally find that the existing human pose condition [17, 44] and temporal embeddings are insufficient to fully capture the expression of human motion. On the one hand, while the human pose represents the spatial information of a template body, it fails to capture local details (e.g., garment deformation). On the other hand, existing temporal embeddings [33, 51] maintain temporal continuity in rendering by sacrificing geometric awareness of the human body, limiting their effectiveness for animation. In this paper, we aim to strike an ideal balance in both rendering and animating of human avatars by introducing a hierarchical spatio-temporal context. To this end, we refine the input conditions for non-rigid warping by modeling a motion sequence, ensuring correspondence of Gaussians across different observations.
The desired motion sequence should encapsulate rich spatio-temporal information to effectively guide the non-rigid warping process. Previous NeRF[38]-based methods [5] construct the motion sequence with human pose residuals to represent human motion. However, the human pose, as an inherently global representation, struggles to capture finer motion details. The explicit structure of 3DGS enables localized temporal modeling, facilitating the construction of a hierarchical motion sequence by encoding motion residuals from template vertices in SMPL(-X). This coarse-to-fine, pose-to-vertex design (as illustrated in Fig. 1) for the motion sequence captures hierarchical spatial motion while guaranteeing temporal continuity. How to determine the optimal stride for motion sequence remains a significant challenge. The motion sequence with a larger stride captures longer motion but loses details between adjacent frames, whereas the motion sequence with a smaller stride covers a shorter motion range but retains richer inter-frame details. To further incorporate detailed human motion, we propose a multi-scale sampling strategy for the motion sequence. The final input to the non-rigid network consists of motion sequences sampled at multiple temporal scales, arranged in a coarse-to-fine manner. This hierarchical temporal design combines both motion range and inter-frame details, further enhancing the model’s generalization ability in rendering from novel views.
We test our pipeline, GAST, on I3D-Human dataset [5] and DNA-Rendering dataset [6], which contain complex garments motions beyond the scope of the template body. Our method outperforms other baselines on these datasets, achieving compatibility in both rendering and animating. To summarize, our contributions are three-fold as follows:
  • 1)
    GAST, the first framework for sequential Gaussian avatars, achieving a balance between high-fidelity rendering and animation through spatio-temporal context;
  • 2)
    a hierarchical motion sequence that combines global human pose and localized vertex residuals, guiding non-rigid warping from coarse to fine;
  • 3)
    a multi-scale sampling strategy for selecting strides for motion sequences, integrating both broader motion dynamics and finer inter-frame details.
2 Related Work
2.1 Human Avatars Based on SMPL(-X)
Recent advances in digital human modeling have benefited greatly from the SMPL(-X) series [37, 42], which represents the human body by decomposing it into distinct shape components and pose-related motions, utilizing 3D mesh scanning techniques and principal component analysis (PCA). The pose blend shapes in SMPL(-X) enable human body deformation through a linear blend of poses of different joints, significantly improving the efficiency of animating digital humans. The widespread adoption of SMPL(-X) has further been boosted for human body animation, driven by methods [7, 46] that estimate parameters of SMPL(-X) from 2D image inputs. Despite its wide applications, the SMPL(-X) framework has its limitations. The skin-tight 3D meshes, based on pose blend shapes, struggle to model bodies with garments or accessories. This brings a challenge for complex human rendering based on SMPL(-X), particularly when dealing with garments and accessories.
2.2 SMPL(-X) Poses Guided Neural Human
Implicit neural methods, such as NeRF [38] and 3DGS [23], have become key contributors to digital human reconstruction due to their high-quality rendering capabilities. Some methods get inspiration from SMPL(-X) to build neural animatable 3D human avatars. These approaches typically map the T-pose to a canonical space and apply LBS weights to perform rigid transformations. Animation is then realized by mapping points from the canonical space to the observation space, with each frame’s human pose determining the transformation. NeRF-based approaches [49, 43, 25, 5, 14, 2, 4, 3, 9, 12, 13] focus on reconstructing human avatars from monocular or multi-view synchronized videos. For more accurate human geometry, [47] incorporate smooth priors from neural Signed Distance Functions (SDF). More recent advances [32, 35, 55, 28, 17, 44, 27, 18, 40, 22, 16, 29, 34, 53, 24, 39, 21, 54] leverage 3DGS, resulting in faster and high-fidelity human rendering. We also notice that some works [5, 18] model a human pose sequence as the context to improve rendering quality. But they suffer from low-speed rendering and lack fine details of motion caused by simply integrating coarse human pose embedding with NeRF.
2.3 Human Rendering with Temporal Embeddings
The earliest modeling of dynamic radiance field [11, 30, 31, 50, 8, 1, 36, 41, 20, 10, 45] can be traced back to temporal embeddings in general scenes, where each frame’s observation is obtained by constructing a canonical space and a time-conditioned deformation field. Motivated by these, a stream of human researches [33, 51] focus on pure rendering quality, employing temporal embeddings instead of human pose to encode each frame of human videos. These methods often achieve high-quality, temporally continuous rendering. However, they struggle with the lack of geometric constraints from human pose, making parameterized animating unfeasible.
3 Preliminary
SMPL(-X) Series [37, 42] adopt a parameterized framework to represent human bodies across diverse shapes and poses. For each frame, the human mesh is derived by deforming a canonical template mesh based on shape and pose parameters. Specifically, a 3D point $\mathbf{x}$ on the canonical mesh is warped to the corresponding point on the deformed mesh in the observation space, following
$\mathbf{LBS}(\mathbf{x},{\mathbf{B}_{k}},\mathbf{\omega}_{k}(\mathbf{x}))=\sum_{k=1}^{K}\mathbf{\omega}_{k}(\mathbf{x})\mathbf{B}_{k}\mathbf{x},$ (1)
where $K$ is the total bone number and $\mathbf{B}_{k}$ is the transformation matrix for each bone. Specifically, $\mathbf{B}_{k}$ consists of a rotation and translation matrix $\mathbf{R}_{k}$ and $\mathbf{b}_{k}$. Here, $\mathbf{R}_{k}$ represents the global rotation of each joint, influenced by the local joint rotations $\bm{r}$, while $\mathbf{b}_{k}$ denotes each joint’s translation, determined by the joint positions $\bm{j}$ and the human shape parameters $\beta$. Note that the linear blending weight $\omega^{k}$ depends on $\mathbf{x}$ and is regressed from comprehensive human meshes.
3D Gaussian Splatting [23] utilizes a set of point-based Gaussian primitives to explicitly represent scenes, enabling real-time and high-quality rendering. Each Gaussian primitive is defined by its center position $\mathbf{x}\in\mathbb{R}^{3}$ and covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$.
To ensure positive semi-definiteness and simplify optimization, the covariance matrix $\Sigma$ is decomposed into a rotation matrix $\mathbf{R}$ and a scaling matrix $\mathbf{S}$ following
$\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T},$ (2)
where $\mathbf{R}$ and $\mathbf{S}$ are derived by a scaling vector $\mathbf{s}\in\mathbb{R}^{3}$ and a quaternion vector $\mathbf{r}\in\mathbb{R}^{4}$ in practice. Additionally, each Gaussian primitive is also assigned a color feature $sh\in\mathbb{R}^{k}$ represented by spherical harmonics (SH) and an opacity $\mathbf{\alpha}$ for rendering. During the rendering process, a splatting technique [23] is used to project each Gaussian primitive onto the 2D image space with a viewing transform $\mathbf{W}$ and the Jacobian $\mathbf{J}$ of the projective transformation’s affine approximation. The transformed covariance $\mathbf{\Sigma}^{\prime}$ in the camera’s coordinates is
$\mathbf{\Sigma}^{\prime}=\mathbf{J}\mathbf{W}\mathbf{\Sigma}\mathbf{W}^{T}\mathbf{J}^{T}.$ (3)
After projection, the pixel color $C$ is computed by blending $N$ ordered Gaussian primitives overlapping at the pixel as
$C=\sum_{i\in N}c_{i}\alpha^{\prime}_{i}\prod_{j=1}^{i-1}(1-\alpha^{\prime}_{j}),$ (4)
where $c_{i}$ is computed from SH feature $sh$ and $\alpha^{\prime}_{i}$ is the product of $\alpha_{i}$ and probability density of $i$-th 2D Gaussian.
For optimization, a photometric loss is defined by a combination of $L_{1}$ and SSIM [48] losses:
$L_{photo}=\lambda_{1}L_{1}(\hat{I},I)+(1-\lambda_{1})(1-SSIM(\hat{I},I)),$ (5)
where $\hat{I}$ and $I$ denote the rendered and ground-truth images, and $\lambda_{1}$ controls the balance between the two terms.
4 Method
Fig. 2 shows the framework of our method. We use the explicit point-based 3DGS as the representation of the human body. Given a collection of input cameras and images, we optimize a set of Gaussian primitives $\{\mathcal{G}_{i}\}_{i=1}^{i=n}$ to fit the body’s shape and appearance. Each Gaussian primitive $\mathcal{G}_{i}$ includes the center position $\mathbf{x}$, scaling vector $\mathbf{s}$, rotation quaternion vector $\mathbf{r}$, color feature $sh$ and opacity $\mathbf{\alpha}$, where $\mathbf{x}$ is initialized from the SMPL template vertices $\{\mathbf{T}_{i}\}_{i=1}^{i=N}$. During the optimization process, we first apply non-rigid deformation with hierarchical spatio-temporal conditions to capture more accurate and fine-grained details (Sec. 4.1). Next, we adopt Linear Blend Skinning (LBS) to map canonical Gaussian primitives $\mathcal{G}$ to observation space and render images via differentiable splitting (Sec.4.3).
Figure 2: Overview of GAST. We first initialize canonical Gaussian positions $\mathbf{x}$ with SMPL template vertexes. We propose a hierarchical spatio-temporal context for non-rigid deformation to better capture complex human motion. The skeleton motion condition $f_{\Delta\mathcal{P}}$ aims to distinguish the appearance variation caused by overall body movement. The point motion condition $f_{\mathcal{V}}$, sampled from the vertex motion template $\mathcal{F}_{\mathcal{V}}$ (points in different colors represent different motions), is designed to capture finer-grained temporal information in local regions. We utilize an MLP $\mathcal{E}_{non-rigid}$ to predict the non-rigid deformation of each Gaussian based on both $f_{\Delta\mathcal{P}}$ and $f_{\mathcal{V}}$ conditions. The non-rigid deformed Gaussians are then warped into observation space via the standard LBS transformation for rendering.
4.1 Non-Rigid Deformation with Hierarchical Spatio-temporal Context
Coarse Skeleton Motion. As illustrated in Fig. 1, we introduce a coarse-to-fine spatio-temporal context to excavate additional conditions for non-rigid deformation. Typically, most human 3D avatar methods use the current frame pose as a condition to predict non-rigid deformation [17, 44]. While this provides spatial information to distinguish the motion of different body parts, it fails to capture temporal motion changes, as discussed in [5]. Therefore, given a frame at time $t$, we consider a sequence of regularly interval-sampled frames $\mathcal{T}$ to model inter-frame body motion variations, defined as
$\mathcal{T}=\{t-s,t-2s,...,t-Ls\},$ (6)
where $L$ is the sequence length and $s$ is the time interval. The coarse motion of body skeletons at each time step $t\in\mathcal{T}$ can be derived by calculating the difference of poses between adjacent frames (as illustrated in Fig. 1), following
$\Delta\mathcal{P}=\{\Delta P^{t}=\delta(P^{t},P^{t-s})|t\in\mathcal{T}\},$ (7)
where $P\in\mathbb{R}^{K\times 3}$ is the body pose and $\delta$ denotes the difference $\Delta\mathcal{P}$ between two poses in axis-angle form. We employ an MLP $\mathcal{E}_{\Delta\mathcal{P}}$ to encode the sequential skeleton motion $\Delta\mathcal{P}\in\mathbb{R}^{L\times K\times 3}$, flattened along the temporal dimension, into a skeleton motion embedding $f_{\Delta\mathcal{P}}\in\mathbb{R}^{32}$, which serves as a condition for the subsequent non-rigid deformation. Further details are provided in the supplementary material.
Fine Vertex Motion. Compared to previous implicit NeRF-based methods, the point-based 3DGS representation enables us to explore more fine-grained temporal motion information. We derive a point-wise velocity vector $v_{i}$ for each Gaussian primitive $\mathcal{G}_{i}$ to model the fine-grained localized body motion, which is beyond the scope of pose skeleton motion $\Delta\mathcal{P}$. By measuring the position variations across adjacent time steps, we calculate per-frame velocity
$v^{t}=\frac{\mathbf{x_{o}}^{t}-\mathbf{x_{o}}^{t-s}}{s},$ (8)
where $\mathbf{x_{o}}$ denotes the warped coordinates in observation space. However, since the positions of Gaussian primitives are updated during optimization, it is not stable to obtain $v_{i}$ based on such a dynamic variable. Moreover, transforming the canonical Gaussians into the observation space requires non-rigid deformation first, which leads to a circular dependency and conflicts with the current velocity design. To this end, we introduce a motion template field $\mathcal{F}_{\mathbf{V}}=\{\mathbf{V}_{i}\}_{i=1}^{i=N}$ that stores the velocity of each SMPL vertex, and $v_{i}$ for each Gaussian primitive is sampled from $\mathcal{F}_{\mathbf{V}}$ based on the distance between query points and vertexes in this template. Specifically, given a time step $t$ and the corresponding body pose $P^{t}=\{(\mathbf{R}_{k}^{t},\mathbf{b}_{k}^{t})|k\in K\}$, the SMPL template vertexes in $\mathbf{T}$ can be warped into observation space with the template skinning weights $\mathbf{W}$ and standard linear blend skinning, following
$\mathbf{T_{o}}^{t}=\mathbf{LBS}(\mathbf{T},\mathbf{B}^{t},\mathbf{W}),$ (9)
where $\mathbf{B}^{t}$ is the transformation matrix derived from rotation and translation matrix $\mathbf{R}^{t}$ and $\mathbf{b}^{t}$. $\mathbf{W}$ is the template SMPL LBS weights. The velocity $\mathbf{V}$ of each template vertexes $\mathbf{T}$ can be further derived as
$\mathbf{V}^{t}=\frac{\mathbf{T_{o}}^{t}-\mathbf{T_{o}}^{t-s}}{s}.$ (10)
We sample the velocity from this motion template field $\mathcal{F}_{\mathbf{V}}$ for each Gaussian primitive. To obtain robust sampled velocities and model body’s local region information, we input the $\tau$ nearest template vertexes’ velocities into an MLP $\mathcal{E}_{knn}$ to learn a local velocity embedding
$e_{i}^{t}=\mathcal{E}_{knn}(\{\mathbf{V}_{j}^{t}\}),\quad j\in\mathbf{KNN}(\mathbf{T},\mathbf{x}_{i}),$ (11)
where $\mathbf{KNN}(\mathbf{T},\mathbf{x}_{i})$ denotes the $\tau$ nearest SMPL template vertexes of the canonical Gaussian position $\mathbf{x}_{i}$. Then we apply an MLP $\mathcal{E}_{\mathcal{V}}$ to encode each Gaussian primitive’s sequential motion state $\mathcal{V}=\{e^{t}|t\in\mathcal{T}\}$ into a point-wise sequential condition $f_{\mathcal{V}}\in\mathbb{R}^{64}$ for non-rigid deformation.
Non-Rigid Deformation. Given the coarse skeleton motion $f_{\Delta\mathcal{P}}=\mathcal{E}_{\Delta\mathcal{P}}(\Delta\mathcal{P})$ condition and fine point-wise velocity condition $f_{\mathcal{V}}=\mathcal{E}_{v}(\mathcal{V})$, we utilize an MLP to predict each Gaussian’s non-rigid deformation as
$\delta\mathbf{x},\delta\mathbf{s},\delta\mathbf{r}=\mathcal{E}_{non-rigid}(\mathbf{x},P,f_{\Delta\mathcal{P}},f_{v}).$ (12)
The deformed canonical Gaussian $\mathcal{G}^{\prime}$ is
$\mathbf{x^{\prime}}$ $=\mathbf{x}+\delta\mathbf{x},$ (13)
$\mathbf{s^{\prime}}$ $=\mathbf{s}+\delta\mathbf{s},$ (14)
$\mathbf{r^{\prime}}$ $=\mathbf{r}\cdot\delta\mathbf{r},$ (15)
where $\cdot$ denotes the multiplication of two quaternions.
4.2 Multi-scale Sequence Sampling
To capture both the overall motion trend and the inter-frame motion details, we adopt a multi-scale sequence sampling strategy. Specifically, we sample a series of motion sequences at several progressively increasing intervals to get
$\mathcal{S}=\{s=s_{0}+i\Delta s\}_{i=0}^{i=m},$ (16)
where $\Delta s$ is the increasing rate of sampling interval $s$. We then input the multi-scale sampled sequence motions into the encoder $\mathcal{E}_{\Delta\mathcal{P}}$ and $\mathcal{E}_{\mathcal{V}}$ to obtain hierarchical temporal motion embeddings
$f_{\Delta\mathcal{P}}$ $=\mathcal{E}_{\Delta\mathcal{P}}(\{\Delta\mathcal{P}_{s}\}),$ (17)
$f_{\mathcal{V}}$ $=\mathcal{E}_{\mathcal{V}}(\{\mathcal{V}_{s}\}),$ (18)
where $s\in\mathcal{S}$. In practice, we concatenate all skeleton motion conditions, $\Delta\mathcal{P}_{s}$ for $s\in\mathcal{S}$, and localized vertex motion conditions, $\mathcal{V}_{s}$ for $s\in\mathcal{S}$, across different sampling scales. These are then input to $\mathcal{E}_{\Delta\mathcal{P}}$ and $\mathcal{E}_{\mathcal{V}}$, respectively. More details are shown in the supplementary material.
4.3 Optimization
Rigid Deformation. We utilize the standard $\mathbf{LBS}$ operation to map the non-rigid deformed Gaussians $\mathcal{G}_{o}$ into the observation space, following
$\mathbf{x_{o}}$ $=\mathbf{LBS}(\mathbf{x^{\prime}},\mathbf{B},\mathbf{\omega}),$ (19)
$\mathbf{R_{o}}$ $=\sum_{k=1}^{K}\mathbf{\omega}_{k}(\mathbf{x^{\prime}})\mathbf{B}_{k}\mathbf{R}^{\prime},$ (20)
where $\mathbf{R}^{\prime}$ is the rotation matrix derived from non-rigid deformed Gaussian’s rotation quaternion $\mathbf{r}^{\prime}$ and $\mathbf{R_{o}}$ represents the rotation matrix in observation space. Following the previous method [17], we utilize an MLP $\mathcal{E}_{lbs}$ to predict the LBS weight offsets for each query canonical point and update the sampled weights from the nearest SMPL vertex as
$\mathbf{\omega}_{k}(\mathbf{x})=\mathbf{\omega}_{k}^{SMPL}(\mathbf{x})+\mathcal{E}_{lbs}(\mathbf{x}).$ (21)
Similar to [49, 5, 17], we introduce a pose refinement MLP $\mathcal{E}_{pose}$ to refine the pose estimate from SMPL for a better fit to the human body.
Loss Function. With the transformed Gaussian primitives $\mathcal{G}_{o}$ in observation space, we apply the standard splitting [23] to render images as
$I=\mathbf{Splatting}(\mathbf{x_{o}},\mathbf{R_{o}},s^{\prime},\alpha,sh).$ (22)
During the optimization process, we employ four loss functions as supervision, summarized as
$\mathcal{L}=\lambda_{1}\mathcal{L}_{color}+\lambda_{2}\mathcal{L}_{ssim}+\lambda_{3}\mathcal{L}_{lpips}+\mathcal{L}_{mask},$ (23)
where $\lambda$ is used to balance the weight of different losses. $\mathcal{L}_{mask}$ [17] is an $L_{2}$ loss between the rendered $\alpha$ and human body mask. Additional details of the pipeline are provided in the supplementary material.
5 Experiments
5.1 Datasets
DNA-Rendering [6] is a challenging realistic human-centric rendering dataset. It contains diverse scenes from everyday life to professional occasions. We choose 6 typical sequences with loose-fitting garments (1_0206_04, 2_0007_04, 2_0019_10, 2_0044_11, 2_0051_09, 2_0813_05) for experiments. We adopt 24 views for training and 6 views for testing. Novel view rendering results compared with other baselines are reported.
I3D-Human [5] comprises multi-view video frames of humans with loose clothing and complex movements. Compared to previous datasets such as ZJU-MoCap [43], which are mainly collected under controlled speeds and tight-fitting garments, I3D-Human contains scenes closer to daily life. We conduct experiments on 4 sequences following the data split in Dyco [5] and all training and testing data are retained to be the same across all comparison methods. For each sequence, we use 4-5 views for training and the rest for testing.
Table 1: Quantitative Results on DNA-Rendering Dataset. We report the performance of novel view rendering. Our GAST outperforms previous state-of-the-art human modeling methods. We mark the best and the second best methods in cells. LPIPS*=LPIPS ×$\times 10^{3}$.
Subject: 1   0206   04 2   0007   04 2   0019   10 2   0044   11 2   0051   09 2   0813   05
[…]
3DGS-Avatar [44] 26.68 0.9434 48.31 28.10 0.9487 53.86 32.09 0.9705 28.87 28.28 0.9506 35.68 25.06 0.9530 41.39 31.75 0.9692 29.94
GART [27] 27.19 0.9454 56.47 28.22 0.9525 56.46 31.13 0.9686 35.16 28.90 0.9561 41.27 26.47 0.9628 42.21 32.06 0.9729 35.74
GauHuman [17] 28.49 0.9534 46.45 28.32 0.9501 52.16 32.55 0.9716 27.91 31.02 0.9670 28.85 26.24 0.9610 38.21 33.79 0.9776 25.14
GAST(Ours) 30.81 0.9649 39.47 29.63 0.9566 43.72 34.97 0.9777 23.53 32.62 0.9747 23.83 28.45 0.9685 34.06 35.81 0.9843 20.83
Figure 3: Qualitative Results on DNA-Rendering Dataset [6]. We compare the visual quality of novel view rendering with previous SOTA methods. To present the differences more clearly, we zoom into the local region and compute the error maps compared with ground truth images. The results show that our method achieves competitive results on both the overall and local region qualities.
Table 2: Quantitative Results on I3D-Human Dataset. We report the performance of novel view rendering. The proposed method achieves competitive results compared with Dyco [5], GauHuman [17], and 3DGS-Avatar [44].
Subject: ID1   1 ID1   2 ID2   1 ID3   1
[…]
3DGS-Avatar [44] 29.78 0.9600 31.40 31.29 0.9627 28.09 29.63 0.9604 37.33 32.73 0.9602 39.47
Dyco [5] 31.53 0.9648 25.17 31.76 0.9646 25.60 29.95 0.9651 29.46 32.78 0.9631 33.06
GauHuman [17] 29.47 0.9571 39.46 30.42 0.9585 38.87 28.51 0.9546 49.46 32.10 0.9546 53.67
GAST(Ours) 31.81 0.9652 28.64 31.88 0.9636 30.01 31.35 0.9669 34.04 33.47 0.9633 37.90
Figure 4: Qualitative Results on I3D-Human Dataset [5]. We compare the visual quality of novel view rendering with SOTA methods.
5.2 Baselines and Metrics
We compare against the state-of-the-art human modeling methods including Dyco [5], 3DGS-Avatar [44], GART [27], and GauHuman [17]. Dyco is an implicit NeRF-based method that adopts pose sequence information for temporal modeling. We compare performance with Dyco to showcase the better efficacy of our point-based fine-grained spatio-temporal context design. 3DGS-Avatar, GART, and GauHuman are explicit 3DGS-based methods, and all of them are originally designed for monocular inputs. For fairness, we extend them to multi-view input under the same settings. We compare these methods to demonstrate that finer rendering details can be achieved with our hierarchical sequential conditions for Gaussian points.
We report three key metrics: peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM) [48], and learned perceptual image patch similarity (LPIPS) [52].
5.3 Comparison
Figure 5: Qualitative Ablation Results. We compare the visual influence of different components on ID2_1 scene of I3D-Human dataset.
Comparison on DNA-Rendering. Tab. 1 shows the quantitative results on DNA-Rendering dataset. The proposed GAST outperforms previous state-of-the-art human modeling methods across all metrics. To demonstrate the visual improvement, we compare the quality of rendering images in Fig. 3. Previous methods, which lack the temporal motion information for non-rigid deformations, are unable to predict the movement of Gaussian primitives properly, resulting in blurred renderings in regions with complex textures (scene 1_0206_04 in Fig. 3). Besides, the results in Fig. 3 show that the compared methods fail to capture the proper shape variation caused by movement (see the white dress in scene 2_0051_09). Although these methods achieve a smooth appearance in these areas, the error maps show that they fail to closely match the actual shape. We leverage the spatio-temporal context to predict Gaussian deformations, allowing our method to capture more accurate detailed appearance variations caused by human motions.
Comparison on I3D-Human. The quantitative results in Tab. 2 show that our method surpasses the previous SOTA method on most metrics, demonstrating our competitive performance on the I3D-Human dataset. Previous NeRF-based Dyco [5] utilizes pose variation as a condition to model human motions. However, it is limited to capturing only overall body motion trends and lacks the capacity for finer-grained modeling in local regions. For example, in ID1_1 and ID2_1 shown in Fig. 4, Dyco fails to model the information in regions far from the human body skeleton, leading to rendering results that do not accurately reflect the true motion. Our method designs more fine-grained conditions for the explicit point-based Gaussian representation to excavate both spatial and temporal motion information hierarchically. The rendering details in Fig. 4 demonstrate the improvement benefiting from such design in these local regions. More results are shown in supplementary materials.
Table 3: Quantitative Ablation Results on I3D-Human Dataset. The Vanilla Non-Rigid MLP in (b) only utilizes Gaussian positions $\mathbf{x}$ and the current frame pose $P$ as the conditions.
Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS* $\downarrow$
(a) Baseline 29.88 0.9570 38.12
(b) (a)+Vanilla Non-Rigid MLP. 31.18 0.9614 35.78
(c) (b)+$\Delta\mathcal{P}$ Conditions. 31.89 0.9640 33.86
(d) (c)+ $\mathcal{V}$ Conditions. 32.05 0.9646 32.82
(e) (d)+Multi-scale Sampling. (Ours) 32.13 0.9647 32.65
5.4 Ablation Study
In this section, we ablate the influence of various components in our methods on I3D-Human dataset. Tab. 3 shows the average metrics of 4 sequences under different settings.
Skeleton motion condition $\Delta\mathcal{P}$ denotes the overall human body movement. Compared to predicting non-rigid deformation only with the Gaussian position, incorporating this additional global temporal information allows our method to better model the overall human shape and achieve higher rendering quality, as demonstrated in Tab. 3 (c).
Fine-grained point-wise motion condition $\mathcal{V}$ describes the local region’s variation in a more detailed manner. Such point-based motion information captures the relationship between motion and appearance variations in areas beyond the human skeleton. Tab. 3 (d) and Fig. 5 (d) show that our method achieves higher rendering details in local regions.
Multi-scale sequence sampling is designed to capture both the overall motion trend and the fine-grained inter-frame motion details. With temporal conditions sampled at different intervals, our method is able to predict each Gaussian’s deformation more robustly, leading to better rendering quality as shown in Tab. 3 (e) and Fig. 5 (e).
6 Limitations and Conclusion
Limitations. Compared to prior NeRF-based approaches, our method models the human body with the point-based Gaussian representation, which may introduce needle-like artifacts during rendering. As indicated in Tab. 2, the metric on LPIPS is slightly inferior to that of NeRF-based Dyco [5]. We notice that some subsequent works [19, 15] on 3DGS optimize the artifacts by introducing surface-related smooth regularization. Generative model-based human avatar [26] could also potentially be introduced as a guarantee of high-fidelity rendering. However, these are beyond the scope of our motion modeling and will serve as future work to improve rendering quality.
Conclusion. In this paper, we propose GAST, a 3DGS-based framework that integrates hierarchical spatial-temporal information for 3D human modeling. Specifically, we introduce a coarse-to-fine motion context for the non-rigid deformation of Gaussian primitives. At the spatial dimension, the global and local region transformations for Gaussians are learned based on the overall inter-frame human pose variations and point-wise motions. To capture stable temporal motion information for each point, we introduce an SMPL vertex velocity template and sample the velocity of each Gaussian point from it. Along the temporal dimension, we apply a multi-scale sampling strategy to capture both long-term motion trends and inter-frame details, supporting more robust predictions of Gaussian deformations. Through the above design, we have achieved state-of-the-art rendering of human avatars with complex garments and motions.
Supplementary Material
7 Details of Skeleton Motion $\Delta P$
The pose $P$ is represented as the rotations of all joints in axis-angle form:
$P=\{\theta_{1},\theta_{2},...,\theta_{K}\},$ (24)
where $\theta_{j}\in\mathbb{R}^{3}$ describes the $j$-th joint’s relative rotation with its parent in axis-angle form [37]. We follow Dyco [5] to derive each joint’s rotation variation $\Delta\theta$ between adjacent frames as its skeleton motion representation. Specifically, the rotation of $j$-th joint in matrix form can be obtained according to Rodrigues formula [37]:
$\mathbf{R}_{j}=Rodrigues(\theta_{j}),$ (25)
where $\mathbf{R}_{j}$ is a 3 $\times$ 3 rotation matrix. Given 2 consecutive poses $P^{t}$ and $P^{t-s}$, the temporal variation of $j$-th joint’s rotation is computed as a relative transformation matrix $\Delta\mathbf{R}_{j}^{t}$:
$\Delta\mathbf{R}_{j}^{t}=\mathbf{R}_{j}^{t}{\mathbf{R}_{j}^{t-s}}^{-1}.$ (26)
Then we convert the 3 $\times$ 3 matrix $\Delta\mathbf{R}_{j}^{t}$ back to the axis-angle form $\Delta\theta_{j}^{t}\in\mathbb{R}^{3}$. Now, the skeleton motion at time $t$ is
$\Delta P^{t}=\{\Delta\theta_{1}^{t},\Delta\theta_{2}^{t},...,\Delta\theta_{K}^{t}\},$ (27)
which is simplified as $\Delta P^{t}=\delta(P^{t},P^{t-s})$ in Eq. (7).
8 Datasets
I3D-Human Dataset [5] We use images with resolution 512 $\times$ 512 for experiments following Dyco [5].
DNA-Rendering Dataset [6] We use the original images’ resolution 512 $\times$ 612 for experiments. For the dataset split, we sample 24 evenly distributed views from camera IDs 0–48 for training and 6 additional uniformly distributed views from IDs 48–60 for novel view testing.
9 Motion Embeddings for Non-Rigid MLP
The embeddings $f_{\Delta\mathcal{P}}$ and $f_{\mathcal{V}}$ in Eq. (12) describe the skeleton and local region motions, respectively. For each frame, all Gaussian primitives $\{\mathcal{G}_{i}\}$ share the same skeleton motion embedding $f_{\Delta\mathcal{P}}\in\mathbb{R}^{32}$ and we concat it with $\mathcal{G}_{i}$’s point-wise motion embedding ${f_{\mathcal{V}}}_{i}\in\mathbb{R}^{64}$ as a new condition $f_{i}^{\prime}\in\mathbb{R}^{96}$. $f_{i}^{\prime}$ is then input to a 2-layer MLP $\mathcal{E}_{non-rigid}$ to predict the corresponding non-rigid deformation.
10 Details of Loss Functions
$\mathcal{L}_{color}$ is the $L_{1}$ loss between the rendered image $I$ and the ground truth $I_{gt}$:
$\mathcal{L}_{color}=|I-I_{gt}|.$ (28)
We use $\mathcal{L}_{ssim}$ to constrain the structure similarity between rendered image and the ground truth, which is given by
$\mathcal{L}_{ssim}=1-\text{SSIM}(I,I_{gt}),$ (29)
where $\text{SSIM}(\cdot)$ is the SSIM metric. Additionally, we use the LPIPS loss to ensure the perceptual similarity:
$\mathcal{L}_{lpips}=\text{LPIPS}(I,I_{gt}),$ (30)
where $\text{LPIPS}(\cdot)$ is the LPIPS metric. Following GauHuman [17] and 3DGS-Avatar [44], we employ a mask loss to ensure that Gaussian primitives are accurately localized within their designated regions:
$\mathcal{L}_{mask}={||M-M_{gt}||}_{2}$ (31)
where $M_{gt}$ is the foreground mask and $M$ is the accumulated $\alpha$ value:
$M=\sum\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).$ (32)
Figure 6: Additional Novel View Qualitative Results on DNA-Rendering Dataset [6]. We provide both the complete image and localized error map comparisons against other methods for a more comprehensive visualization.
11 Details of Optimization
In our experiments, the number $\lvert\mathcal{S}\rvert$ (in Eq. (16)) of sequence sampled in different scales is set to 3 and each sequence length $L$ (in Eq. (6)) is set to 8. We use the $\tau=8$ nearest SMPL template vertexes’ velocities to get each Gaussian primitive’s local velocity embedding $e_{i}$ (in Eq. (11)). Additionally, we utilize the adaptive densification in  [23] to control the number of Gaussian primitives.
For DNA-Rendering dataset, we use the SMPL-X [42] model with pose dimension $P\in\mathbb{R}^{25\times 3}$ in our experiments. We adopt the SMPL-X template with $N=10475$ vertexes as Gaussian initialization. We optimize 30k iterations and set the loss weights $\lambda_{0}=1.0,\lambda_{1}=0.01,\lambda_{2}=0.01$. The initial sampling step and the increasing interval are $s_{0}=1$ and $\Delta s=2$, respectively. The final sample scales are $\mathcal{S}=\{1,3,5\}$ (in Eq. (16)). For I3D-Human dataset, we follow Dyco [5] to use SMPL [37] model in our experiments for fairness. The number of SMPL template vertexes used for Gaussian initialization is $N=6890$ and the corresponding pose dimension is $P\in\mathbb{R}^{23\times 3}$. We optimize 15k iterations and set the loss weights $\lambda_{0}=1.0,\lambda_{1}=0.1,\lambda_{2}=0.1$. We set the increasing interval $\Delta s=9$. The initial sampling step is $s_{0}=24$ and the final sample scales are $\mathcal{S}=\{24,33,42\}$ (in Eq. (16)). Note that the experiment in Tab. 3 (d) adopts single sampling step that $\mathcal{S}=\{s_{0}\}$.
12 Additional Results
12.1 Comparisons of Novel View Rendering
Figure 7: Additional Novel Pose Qualitative Results on I3D-Human Dataset [5].
Fig. 6 shows additional qualitative comparisons with other SOTA methods on the DNA-Rendering dataset [6]. Our approach excels in regions where motion causes notable appearance variations, producing results that conform to the true shape more accurately, as exemplified by the sleeve (top row) and the skirt hem (bottom row) in Fig. 6.
12.2 Comparisons of Novel Pose Rendering
Table 4: Quantitative Results of Novel Pose Renderings on I3D-Human Dataset.
Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS* $\downarrow$
3DGS-Avatar [44] 29.53 0.9554 38.58
Dyco [5] 30.30 0.9582 33.52
GauHuman [17] 29.31 0.9525 48.29
GAST(Ours) 30.39 0.9582 38.15
Tab. 4 reports the quantitative results of novel pose rendering on the I3D-Human dataset [5]. Our method achieves competitive performance compared with SOTA human modeling methods. Fig. 7 shows the qualitative comparisons of novel pose rendering and we utilize error maps to visualize the differences between various methods more clearly. Compared to Dyco [5], which relies solely on joint motions as conditions, our GAST leverages the motion state of each Gaussian primitive to predict deformations, enabling more flexible modeling of motions in regions distant from the human body. The shape of the clothing in Fig. 7 demonstrates that our method can render the appearance consistent with the novel pose motion, leading to more accurate results.
13 Additional Ablations
Different Sample Steps Tab. 5 shows the quantitative metrics with different sequential sampling steps. The results demonstrate that combining all coarse-to-fine temporal motions as the condition leads to more robust non-rigid deformation and better performance.
Table 5: Impact of Different Sampling Steps on I3D-Human.
Sampling Step $\mathcal{S}$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS* $\downarrow$
$\mathcal{S}=\{24\}$ 32.05 0.96468 32.82
$\mathcal{S}=\{33\}$ 32.04 0.96467 33.06
$\mathcal{S}=\{42\}$ 32.06 0.96472 32.87
$\mathcal{S}=\{24,33,42\}$ 32.13 0.96476 32.65
Different KNN Sampled Vertexes Tab. 6 presents the quantitative results on I3D-Human dataset [5] with different numbers of KNN vertexes when deriving the velocity $e_{i}$ of each Gaussian primitive (Eq. 11). It suggests that considering the motion states of all local regions helps obtain a more robust Gaussian primitive’s velocity embedding, contributing to enhancing performance.
Table 6: Impact of Different KNN Vertex Number $\tau$.
KNN num. $\tau$ $\tau=1$ $\tau=3$ $\tau=6$ $\tau=8$
PSNR $\uparrow$ 32.06 32.07 32.08 32.13

### v2

@@
-GAST: Sequential Gaussian Avatars with Hierarchical Spatio-temporal Context
+Sequential Gaussian Avatars with Hierarchical Motion Context
 Abstract
-3D human avatars, through the use of canonical radiance fields and per-frame observed warping, enable high-fidelity rendering and animating.
-However, existing methods, which rely on either spatial SMPL(-X) poses or temporal embeddings, respectively suffer from coarse rendering quality or limited animation flexibility.
-To address these challenges, we propose GAST, a framework that unifies 3D human modeling with 3DGS by hierarchically integrating both spatial and temporal information.
-Specifically, we design a sequential conditioning framework for the non-rigid warping of the human body, under whose guidance more accurate 3D Gaussians can be obtained in the observation space.
-Moreover, the explicit properties of Gaussians allow us to embed richer sequential information, encompassing both the coarse sequence of human poses and finer per-vertex motion details.
-These sequence conditions are further sampled across different temporal scales, in a coarse-to-fine manner, ensuring unbiased inputs for non-rigid warping.
-Experimental results demonstrate that our method combined with hierarchical spatio-temporal modeling surpasses concurrent baselines, delivering both high-quality rendering and flexible animating capabilities.
+The emergence of neural rendering has significantly advanced the rendering quality of 3D human avatars, with the recently popular 3DGS technique enabling real-time performance.
+However, SMPL-driven 3DGS human avatars still struggle to capture fine appearance details due to the complex mapping from pose to appearance during fitting.
+In this paper, we propose SeqAvatar, which excavates the explicit 3DGS representation to better model human avatars based on a hierarchical motion context.
+Specifically, we utilize a coarse-to-fine motion conditions that incorporate both the overall human skeleton and fine-grained vertex motions for non-rigid deformation.
+To enhance the robustness of the proposed motion conditions, we adopt a spatio-temporal multi-scale sampling strategy to hierarchically integrate more motion clues to model human avatars.
+Extensive experiments demonstrate that our method significantly outperforms 3DGS-based approaches and renders human avatars orders of magnitude faster than the latest NeRF-based models that incorporate temporal context, all while delivering performance that is at least comparable or even superior.
+Project page: https://zezeaaa.github.io/projects/SeqAvatar/
+Figure 1: Illustration of the Hierarchical Motion Context Design.
+We model the motion-dependent appearance variations with both coarse skeleton condition $\Delta P$ and fine-grained point-wise velocity condition $\mathcal{V}$.
+$\Delta P$ describes the overall human skeleton motion, which is derived from the difference between the human poses at adjacent frames.
+$\mathcal{V}$ is the point-wise velocity that indicates finer-grained motion in local regions.
+To achieve robust deformation, we propose Spatio-temporal Multi-scale Sampling, which samples the overall motion trend and inter-frame details via diverse time intervals $s$ for $\Delta P$ and $\mathcal{V}$.
 1 Introduction
-Figure 1: Illustration of the Hierarchical Spatio-temporal Context Design.
-We model the motion-dependent appearance variations with both coarse skeleton condition $\Delta P$ and fine-grained point-wise velocity condition $\mathcal{V}$.
-$\Delta P$ describes the overall human pose motion, which is derived from the difference between the human poses at adjacent frames.
-$\mathcal{V}$ is the point-wise velocity that indicates finer-grained motion in local regions.
-Based on $\Delta P$ and $\mathcal{V}$, we utilize an MLP to predict more accurate non-rigid deformation of each Gaussian primitive.
-To achieve robust deformation, we sample the overall motion trend and inter-frame details via diverse time intervals $s$ for $\Delta P$ and $\mathcal{V}$.
-The rendering results demonstrate that our method excels in regions with complex texture details as well as in areas where the appearance varies dramatically with motions.
-Recent research on digital humans has highlighted the efficiency of 3D Gaussian Splatting [23] (3DGS), demonstrating its capability for high-quality and real-time rendering.
-By defining T-pose human Gaussians in a canonical space and employing specific warping techniques, we can render human avatars under various observations.
-The canonical Gaussian units and warping weights are jointly optimized under supervision from human video input.
-The primary bottleneck in this task stems from the limited modeling of non-rigid warping, requiring a trade-off between rendering quality and animating capability.
-One stream of approaches [17, 44] introduces the prior from SMPL(-X) [37, 42] poses, leveraging human poses to guide motions of each observation and relying on Linear Blend Skinning (LBS) for motion warping.
-Although they are well-suited for pose-driven animation, they often struggle with per-frame non-rigid warping in scenarios with complex garments.
-Other approaches [33, 51] utilize temporal embeddings instead of SMPL poses to model each frame independently, achieving high-quality rendering but making it challenging to animate avatars.
-We experimentally find that the existing human pose condition [17, 44] and temporal embeddings are insufficient to fully capture the expression of human motion.
-On the one hand, while the human pose represents the spatial information of a template body, it fails to capture local details (e.g., garment deformation).
-On the other hand, existing temporal embeddings [33, 51] maintain temporal continuity in rendering by sacrificing geometric awareness of the human body, limiting their effectiveness for animation.
-In this paper, we aim to strike an ideal balance in both rendering and animating of human avatars by introducing a hierarchical spatio-temporal context.
-To this end, we refine the input conditions for non-rigid warping by modeling a motion sequence, ensuring correspondence of Gaussians across different observations.
-The desired motion sequence should encapsulate rich spatio-temporal information to effectively guide the non-rigid warping process.
-Previous NeRF[38]-based methods [5] construct the motion sequence with human pose residuals to represent human motion.
-However, the human pose, as an inherently global representation, struggles to capture finer motion details.
-The explicit structure of 3DGS enables localized temporal modeling, facilitating the construction of a hierarchical motion sequence by encoding motion residuals from template vertices in SMPL(-X).
-This coarse-to-fine, pose-to-vertex design (as illustrated in Fig.
-1) for the motion sequence captures hierarchical spatial motion while guaranteeing temporal continuity.
-How to determine the optimal stride for motion sequence remains a significant challenge.
-The motion sequence with a larger stride captures longer motion but loses details between adjacent frames, whereas the motion sequence with a smaller stride covers a shorter motion range but retains richer inter-frame details.
-To further incorporate detailed human motion, we propose a multi-scale sampling strategy for the motion sequence.
-The final input to the non-rigid network consists of motion sequences sampled at multiple temporal scales, arranged in a coarse-to-fine manner.
-This hierarchical temporal design combines both motion range and inter-frame details, further enhancing the model’s generalization ability in rendering from novel views.
-We test our pipeline, GAST, on I3D-Human dataset [5] and DNA-Rendering dataset [6], which contain complex garments motions beyond the scope of the template body.
-Our method outperforms other baselines on these datasets, achieving compatibility in both rendering and animating.
-To summarize, our contributions are three-fold as follows:
+Recent research on digital humans has highlighted the efficiency of 3D Gaussian Splatting (3DGS) [27], demonstrating its capability for high-quality and real-time rendering.
+By defining T-pose human Gaussians in a canonical space and employing specific warping techniques, we can render human avatars from diverse perspectives and in any pose.
+The canonical Gaussian primitives and warping weights are then jointly optimized under supervision from video inputs.
+Animatable human avatar reconstruction is primarily hindered by limited modeling of non-rigid warping, forcing a trade-off between rendering quality and animation capability.
+Current methods  [22, 56] incorporate SMPL(-X) pose priors  [43, 51] to guide motion for each observation and rely on Linear Blend Skinning (LBS) for warping.
+Although these approaches are effective for pose-driven animation, they often struggle with per-frame non-rigid warping in scenarios involving complex garments.
+We experimentally find that current human pose conditions [22, 56] do not fully capture the complex one-to-many mapping from pose to appearance, particularly for 3DGS-based approaches [27].
+These methods [68, 56] rely on the spatial information of a template body in each frame’s human pose to predict non-rigid deformations.
+However, they often overlook local details, such as garment deformations far from the skeleton, and cannot resolve cases where the same pose corresponds to different appearances during complex motions.
+Furthermore, although previous NeRF-based [44] method [8] has attempted to model motion sequences using human pose residuals, the inherently global nature of the pose sequence limits the ability to capture finer motion details.
+Naive pose sequence modeling does not fully account for the explicit characteristics of 3DGS.
+In this paper, we introduce a hierarchical motion context condition for 3DGS-based human avatar modeling to address the limitations of relying solely on human pose, which provides only limited global skeletal information.
+Our approach improves the ability of 3DGS-based methods to accurately capture the complex relationship between human pose and appearance in challenging scenarios.
+Specifically, we design a coarse-to-fine motion condition that incorporates both overall skeletal movements and fine-grained point-wise motions.
+Leveraging the explicit nature of Gaussian primitives, this condition seamlessly integrates into 3DGS-based methods, enabling more precise predictions of complex non-rigid deformations.
+To further enhance robustness, we propose a spatio-temporal multi-scale sampling strategy for constructing the motion context with a larger receptive field.
+Spatially, we consider the motion states of neighboring points within the local region of each Gaussian primitive to obtain more stable motion embeddings.
+Temporally, we capture human motion patterns across multiple time scales, combining long-term trends with fine-grained inter-frame motion details.
+This improves the model’s generalization to complex human movements.
+To summarize, our key contributions are as follows:
   • 1)
-    GAST, the first framework for sequential Gaussian avatars, achieving a balance between high-fidelity rendering and animation through spatio-temporal context;
+    We propose a novel hierarchical motion condition that integrates coarse-to-fine human motion, combining global skeletal poses with localized vertex residuals to enhance non-rigid deformation prediction.
   • 2)
-    a hierarchical motion sequence that combines global human pose and localized vertex residuals, guiding non-rigid warping from coarse to fine;
+    We introduce a spatiotemporal multiscale sampling strategy that expands the receptive field of hierarchical motion context, improving generalization to complex motions.
   • 3)
-    a multi-scale sampling strategy for selecting strides for motion sequences, integrating both broader motion dynamics and finer inter-frame details.
+    Experiments on the I3D-Human [8], DNA-Rendering [9], and ZJU-MoCap [53] datasets show the effectiveness of the proposed SeqAvatar, which is capable of modeling details of the human body in complex motions.
 2 Related Work
-2.1 Human Avatars Based on SMPL(-X)
-Recent advances in digital human modeling have benefited greatly from the SMPL(-X) series [37, 42], which represents the human body by decomposing it into distinct shape components and pose-related motions, utilizing 3D mesh scanning techniques and principal component analysis (PCA).
-The pose blend shapes in SMPL(-X) enable human body deformation through a linear blend of poses of different joints, significantly improving the efficiency of animating digital humans.
-The widespread adoption of SMPL(-X) has further been boosted for human body animation, driven by methods [7, 46] that estimate parameters of SMPL(-X) from 2D image inputs.
-Despite its wide applications, the SMPL(-X) framework has its limitations.
-The skin-tight 3D meshes, based on pose blend shapes, struggle to model bodies with garments or accessories.
-This brings a challenge for complex human rendering based on SMPL(-X), particularly when dealing with garments and accessories.
-2.2 SMPL(-X) Poses Guided Neural Human
-Implicit neural methods, such as NeRF [38] and 3DGS [23], have become key contributors to digital human reconstruction due to their high-quality rendering capabilities.
-Some methods get inspiration from SMPL(-X) to build neural animatable 3D human avatars.
-These approaches typically map the T-pose to a canonical space and apply LBS weights to perform rigid transformations.
-Animation is then realized by mapping points from the canonical space to the observation space, with each frame’s human pose determining the transformation.
-NeRF-based approaches [49, 43, 25, 5, 14, 2, 4, 3, 9, 12, 13] focus on reconstructing human avatars from monocular or multi-view synchronized videos.
-For more accurate human geometry, [47] incorporate smooth priors from neural Signed Distance Functions (SDF).
-More recent advances [32, 35, 55, 28, 17, 44, 27, 18, 40, 22, 16, 29, 34, 53, 24, 39, 21, 54] leverage 3DGS, resulting in faster and high-fidelity human rendering.
-We also notice that some works [5, 18] model a human pose sequence as the context to improve rendering quality.
-But they suffer from low-speed rendering and lack fine details of motion caused by simply integrating coarse human pose embedding with NeRF.
-2.3 Human Rendering with Temporal Embeddings
-The earliest modeling of dynamic radiance field [11, 30, 31, 50, 8, 1, 36, 41, 20, 10, 45] can be traced back to temporal embeddings in general scenes, where each frame’s observation is obtained by constructing a canonical space and a time-conditioned deformation field.
-Motivated by these, a stream of human researches [33, 51] focus on pure rendering quality, employing temporal embeddings instead of human pose to encode each frame of human videos.
-These methods often achieve high-quality, temporally continuous rendering.
-However, they struggle with the lack of geometric constraints from human pose, making parameterized animating unfeasible.
+Neural Rendering.
+Neural rendering techniques have brought significant progress to human reconstruction and rendering.
+In particular, Neural Radiance Fields (NeRF) [44] introduces an implicit scene representation that models color and density using multilayer perceptrons, delivering photorealistic rendering results.
+More recently, point-based 3D Gaussian Splatting [27] utilizes 3D Gaussians to represent scenes explicitly, achieving real-time high-quality rendering.
+Building upon NeRF and 3DGS, subsequent works have advanced neural rendering across various dimensions, e.g., enhancing visual fidelity [1, 2], improving sparse-view reconstruction quality [10, 63, 80, 71, 52], enabling pose-free optimization [67, 37, 61, 58], modeling dynamic scenes [55, 13, 34, 49, 69, 74], and accelerating training and inference [18, 46, 12, 60, 4, 39].
+These developments also benefit human avatar modeling, which demands high-quality and efficient rendering.
+SMPL(-X)-Based Neural Human Modeling.
+The SMPL(-X) family [43, 51] provides a parametric representation of the human body by decomposing it into pose-related and shape-related components using 3D mesh scans and principal component analysis (PCA).
+Its pose blend shapes enables body deformation through joint-wise pose blending, offering an efficient and compact way for animation.
+SMPL(-X) has thus become a cornerstone in human body modeling and animation, with many methods [11, 59] estimating its parameters directly from 2D inputs.
+Recent neural human reconstruction methods integrate SMPL(-X) with implicit representations such as NeRF [44] and 3DGS [27] to enable animatable avatars with high rendering quality.
+Different from purely 2D image animation methods [73, 65, 62, 48], these methods typically register the input data to a canonical T-pose space and use linear blend skinning (LBS) to transform 3D points into observation space based on SMPL(-X) poses.
+NeRF-based approaches [68, 53, 30, 8, 20, 5, 7, 6, 14, 17, 19, 64, 75] focus on monocular or multi-view reconstruction.
+More recently, 3DGS-based methods [36, 41, 81, 32, 22, 56, 31, 23, 47, 26, 21, 33, 40, 78, 29, 45, 25, 79, 76] have gained popularity due to their real-time rendering and high fidelity.
+Some works [8, 23] further incorporate pose sequences as temporal context, but they suffer from slow rendering and lack fine details of motion caused by simply integrating coarse human pose embedding with NeRF.
+Human Rendering with Temporal Embeddings The earliest modeling of dynamic radiance field [16, 34, 35, 70, 13, 3, 42, 50, 24, 15, 57] can be traced back to temporal embeddings in general scenes, where each frame’s observation is obtained by constructing a canonical space and a time-conditioned deformation field.
+Motivated by these, a stream of research [38, 72] focuses on pure rendering quality, employing temporal embeddings instead of human pose to encode each frame of human videos.
+Although these methods achieve high-quality, temporally continuous rendering, they struggle with the lack of geometric constraints from human pose, making parameterized animation unfeasible.
 3 Preliminary
@@
 Here, $\mathbf{R}_{k}$ represents the global rotation of each joint, influenced by the local joint rotations $\bm{r}$, while $\mathbf{b}_{k}$ denotes each joint’s translation, determined by the joint positions $\bm{j}$ and the human shape parameters $\beta$.
-Note that the linear blending weight $\omega^{k}$ depends on $\mathbf{x}$ and is regressed from comprehensive human meshes.
+The linear blending weight $\omega^{k}$ depends on $\mathbf{x}$ and is regressed from comprehensive human meshes.
 3D Gaussian Splatting [27] utilizes a set of point-based Gaussian primitives to explicitly represent scenes, enabling real-time and high-quality rendering.
 Each Gaussian primitive is defined by its center position $\mathbf{x}\in\mathbb{R}^{3}$ and covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$.
-To ensure positive semi-definiteness and simplify optimization, the covariance matrix $\Sigma$ is decomposed into a rotation matrix $\mathbf{R}$ and a scaling matrix $\mathbf{S}$ following
+To ensure positive semi-definiteness and simplify optimization, the covariance matrix $\Sigma$ is decomposed into a rotation matrix $\mathbf{R}$ and a scaling matrix $\mathbf{S}$:
 $\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T},$ (2)
@@
 $\mathbf{\Sigma}^{\prime}=\mathbf{J}\mathbf{W}\mathbf{\Sigma}\mathbf{W}^{T}\mathbf{J}^{T}.$ (3)
-After projection, the pixel color $C$ is computed by blending $N$ ordered Gaussian primitives overlapping at the pixel as
+After projection, the pixel color $C$ is computed by blending $N$ ordered Gaussian primitives overlapping at the pixel:
 $C=\sum_{i\in N}c_{i}\alpha^{\prime}_{i}\prod_{j=1}^{i-1}(1-\alpha^{\prime}_{j}),$ (4)
@@
 For optimization, a photometric loss is defined by a combination of $L_{1}$ and SSIM [66] losses:
-$L_{photo}=\lambda_{1}L_{1}(\hat{I},I)+(1-\lambda_{1})(1-SSIM(\hat{I},I)),$ (5)
-where $\hat{I}$ and $I$ denote the rendered and ground-truth images, and $\lambda_{1}$ controls the balance between the two terms.
+$L_{photo}=\lambda L_{1}(\hat{I},I)+(1-\lambda)(1-SSIM(\hat{I},I)),$ (5)
+where $\hat{I}$ and $I$ denote the rendered and ground-truth images, and $\lambda$ controls the balance between the two terms.
 4 Method
@@
 Each Gaussian primitive $\mathcal{G}_{i}$ includes the center position $\mathbf{x}$, scaling vector $\mathbf{s}$, rotation quaternion vector $\mathbf{r}$, color feature $sh$ and opacity $\mathbf{\alpha}$, where $\mathbf{x}$ is initialized from the SMPL template vertices $\{\mathbf{T}_{i}\}_{i=1}^{i=N}$.
-During the optimization process, we first apply non-rigid deformation with hierarchical spatio-temporal conditions to capture more accurate and fine-grained details (Sec.
+During the optimization process, we first apply a coarse-to-fine motion context to capture more accurate and fine-grained details (Sec.
 4.1).
+To mitigate overfitting, we propose Spatio-Temporal Multi-Scale Sampling to obtain more robust point-wise Gaussian motion, which serves as the embedding for non-rigid deformation (Sec.4.2).
 Next, we adopt Linear Blend Skinning (LBS) to map canonical Gaussian primitives $\mathcal{G}$ to observation space and render images via differentiable splitting (Sec.4.3).
-Figure 2: Overview of GAST.
+Figure 2: Overview of the proposed method.
 We first initialize canonical Gaussian positions $\mathbf{x}$ with SMPL template vertexes.
-We propose a hierarchical spatio-temporal context for non-rigid deformation to better capture complex human motion.
-The skeleton motion condition $f_{\Delta\mathcal{P}}$ aims to distinguish the appearance variation caused by overall body movement.
-The point motion condition $f_{\mathcal{V}}$, sampled from the vertex motion template $\mathcal{F}_{\mathcal{V}}$ (points in different colors represent different motions), is designed to capture finer-grained temporal information in local regions.
-We utilize an MLP $\mathcal{E}_{non-rigid}$ to predict the non-rigid deformation of each Gaussian based on both $f_{\Delta\mathcal{P}}$ and $f_{\mathcal{V}}$ conditions.
+For each Gaussian, we derive both coarse skeleton motion condition $f_{\Delta\mathcal{P}}$ and fine-grained vertex motion condition $f_{\mathcal{V}}$ that sampled from the vertex motion template $\mathcal{F}_{\mathcal{V}}$ (points in different colors represent different motions).
+Based on such hierarchical motion information, we utilize an MLP $\mathcal{E}_{non-rigid}$ to better predict each Gaussian’s non-rigid deformation prediction.
 The non-rigid deformed Gaussians are then warped into observation space via the standard LBS transformation for rendering.
-4.1 Non-Rigid Deformation with Hierarchical Spatio-temporal Context
+4.1 Non-Rigid Deformation with Coarse-to-fine Motion Context
 Coarse Skeleton Motion.
 As illustrated in Fig.
-1, we introduce a coarse-to-fine spatio-temporal context to excavate additional conditions for non-rigid deformation.
+1, we introduce a coarse-to-fine motion context to excavate additional conditions for non-rigid deformation.
 Typically, most human 3D avatar methods use the current frame pose as a condition to predict non-rigid deformation [22, 56].
 While this provides spatial information to distinguish the motion of different body parts, it fails to capture temporal motion changes, as discussed in [8].
-Therefore, given a frame at time $t$, we consider a sequence of regularly interval-sampled frames $\mathcal{T}$ to model inter-frame body motion variations, defined as
+Therefore, given a frame at time $t$, we consider a sequence of regularly interval-sampled frames $\mathcal{T}$ to model inter-frame body motion variations:
 $\mathcal{T}=\{t-s,t-2s,...,t-Ls\},$ (6)
 where $L$ is the sequence length and $s$ is the time interval.
-The coarse motion of body skeletons at each time step $t\in\mathcal{T}$ can be derived by calculating the difference of poses between adjacent frames (as illustrated in Fig.
-1), following
+The coarse motion of body skeletons at each time step $t\in\mathcal{T}$ can be derived by calculating the difference of poses between adjacent frames as illustrated in Fig.
+1, following
 $\Delta\mathcal{P}=\{\Delta P^{t}=\delta(P^{t},P^{t-s})|t\in\mathcal{T}\},$ (7)
 where $P\in\mathbb{R}^{K\times 3}$ is the body pose and $\delta$ denotes the difference $\Delta\mathcal{P}$ between two poses in axis-angle form.
-We employ an MLP $\mathcal{E}_{\Delta\mathcal{P}}$ to encode the sequential skeleton motion $\Delta\mathcal{P}\in\mathbb{R}^{L\times K\times 3}$, flattened along the temporal dimension, into a skeleton motion embedding $f_{\Delta\mathcal{P}}\in\mathbb{R}^{32}$, which serves as a condition for the subsequent non-rigid deformation.
-Further details are provided in the supplementary material.
+We employ an MLP $\mathcal{E}_{\Delta\mathcal{P}}$ to encode the sequential skeleton motion $\Delta\mathcal{P}\in\mathbb{R}^{L\times K\times 3}$, flattened along the temporal dimension, into a skeleton motion embedding $f_{\Delta\mathcal{P}}\in\mathbb{R}^{32}$,
+$f_{\Delta\mathcal{P}}=\mathcal{E}_{\Delta\mathcal{P}}(\Delta\mathcal{P}),$ (8)
+which serves as a condition for the subsequent non-rigid deformation. Further details are provided in the supplementary material.
 Fine Vertex Motion.
@@
 $\mathbf{V}^{t}=\frac{\mathbf{T_{o}}^{t}-\mathbf{T_{o}}^{t-s}}{s}.$ (11)
-We sample the velocity from this motion template field $\mathcal{F}_{\mathbf{V}}$ for each Gaussian primitive.
-To obtain robust sampled velocities and model body’s local region information, we input the $\tau$ nearest template vertexes’ velocities into an MLP $\mathcal{E}_{knn}$ to learn a local velocity embedding
+We then sample each Gaussian’s velocity from this motion template field $\mathcal{F}_{\mathbf{V}}$, which is discussed in Sec. 4.2.
+4.2 Spatio-temporal Multi-scale Sampling (STMS)
+To enhance the robustness of human motion conditions, we propose modeling both local region motion information of the human body and motion patterns across different temporal windows to mitigate overfitting and improve generalization by capturing more comprehensive motion dynamics.
+In the spatial dimension, we sample the $\tau$ nearest template to model the body’s local region motion more robustly.
+Specifically, for each canonical Gaussian primitive $\mathcal{G}_{i}$, the $\tau$ nearest vertexes’ velocities are sampled as input to an MLP $\mathcal{E}_{knn}$ to learn a local point-wise motion embedding
 $e_{i}^{t}=\mathcal{E}_{knn}(\{\mathbf{V}_{j}^{t}\}),\quad j\in\mathbf{KNN}(\mathbf{T},\mathbf{x}_{i}),$ (12)
 where $\mathbf{KNN}(\mathbf{T},\mathbf{x}_{i})$ denotes the $\tau$ nearest SMPL template vertexes of the canonical Gaussian position $\mathbf{x}_{i}$.
-Then we apply an MLP $\mathcal{E}_{\mathcal{V}}$ to encode each Gaussian primitive’s sequential motion state $\mathcal{V}=\{e^{t}|t\in\mathcal{T}\}$ into a point-wise sequential condition $f_{\mathcal{V}}\in\mathbb{R}^{64}$ for non-rigid deformation.
+Similar to Eq.
+7, we then apply an MLP $\mathcal{E}_{\mathcal{V}}$ to encode each Gaussian primitive’s sequential motion embedding $\mathcal{V}=\{e^{t}|t\in\mathcal{T}\}$ into a point-wise sequential condition $f_{\mathcal{V}}\in\mathbb{R}^{96}$
+$f_{\mathcal{V}}=\mathcal{E}_{\mathcal{V}}(\mathcal{V}).$ (13)
+In the temporal dimension, to capture both the overall motion trend and the inter-frame motion details, we adopt a multi-scale sequence sampling strategy.
+In detail, we sample a series of sequences at several progressively increasing intervals to get human motion information across different temporal windows
+$\mathcal{S}=\{s=s_{0}+i\Delta s\}_{i=0}^{i=m},$ (14)
+where $\Delta s$ is the increasing rate of sampling interval $s$.
+We then input the multi-scale sampled sequence motions into the skeleton motion encoder $\mathcal{E}_{\Delta\mathcal{P}}$ and the point-wise motion encoder $\mathcal{E}_{\mathcal{V}}$ to obtain hierarchical temporal motion embeddings.
+Therefore, Eq.
+7 and Eq.
+13 can be revised as
+$f_{\Delta\mathcal{P}}$ $=\mathcal{E}_{\Delta\mathcal{P}}(\{\Delta\mathcal{P}_{s}\}),\quad s\in\mathcal{S}$ (15)
+$f_{\mathcal{V}}$ $=\mathcal{E}_{\mathcal{V}}(\{\mathcal{V}_{s}\}),\quad s\in\mathcal{S}.$ (16)
+In practice, we concatenate all skeleton motion conditions, $\Delta\mathcal{P}_{s}$ for $s\in\mathcal{S}$, and localized vertex motion conditions, $\mathcal{V}_{s}$ for $s\in\mathcal{S}$, across different sampling scales.
+These are then input to $\mathcal{E}_{\Delta\mathcal{P}}$ and $\mathcal{E}_{\mathcal{V}}$, respectively.
+More details are shown in the supplementary material.
 Non-Rigid Deformation.
-Given the coarse skeleton motion $f_{\Delta\mathcal{P}}=\mathcal{E}_{\Delta\mathcal{P}}(\Delta\mathcal{P})$ condition and fine point-wise velocity condition $f_{\mathcal{V}}=\mathcal{E}_{v}(\mathcal{V})$, we utilize an MLP to predict each Gaussian’s non-rigid deformation as
+Given the coarse skeleton motion $f_{\Delta\mathcal{P}}=\mathcal{E}_{\Delta\mathcal{P}}(\{\Delta\mathcal{P}_{s}\})$ condition and fine point-wise velocity condition $f_{\mathcal{V}}=\mathcal{E}_{\mathcal{V}}(\{\mathcal{V}_{s}\})$, we utilize an MLP to predict each Gaussian’s non-rigid deformation as
 $\delta\mathbf{x},\delta\mathbf{s},\delta\mathbf{r}=\mathcal{E}_{non-rigid}(\mathbf{x},P,f_{\Delta\mathcal{P}},f_{v}).$ (17)
@@
 where $\cdot$ denotes the multiplication of two quaternions.
-4.2 Multi-scale Sequence Sampling
-To capture both the overall motion trend and the inter-frame motion details, we adopt a multi-scale sequence sampling strategy.
-Specifically, we sample a series of motion sequences at several progressively increasing intervals to get
-$\mathcal{S}=\{s=s_{0}+i\Delta s\}_{i=0}^{i=m},$ (16)
-where $\Delta s$ is the increasing rate of sampling interval $s$.
-We then input the multi-scale sampled sequence motions into the encoder $\mathcal{E}_{\Delta\mathcal{P}}$ and $\mathcal{E}_{\mathcal{V}}$ to obtain hierarchical temporal motion embeddings
-$f_{\Delta\mathcal{P}}$ $=\mathcal{E}_{\Delta\mathcal{P}}(\{\Delta\mathcal{P}_{s}\}),$ (17)
-$f_{\mathcal{V}}$ $=\mathcal{E}_{\mathcal{V}}(\{\mathcal{V}_{s}\}),$ (18)
-where $s\in\mathcal{S}$.
-In practice, we concatenate all skeleton motion conditions, $\Delta\mathcal{P}_{s}$ for $s\in\mathcal{S}$, and localized vertex motion conditions, $\mathcal{V}_{s}$ for $s\in\mathcal{S}$, across different sampling scales.
-These are then input to $\mathcal{E}_{\Delta\mathcal{P}}$ and $\mathcal{E}_{\mathcal{V}}$, respectively.
-More details are shown in the supplementary material.
 4.3 Optimization
@@
 $\mathbf{R_{o}}$ $=\sum_{k=1}^{K}\mathbf{\omega}_{k}(\mathbf{x^{\prime}})\mathbf{B}_{k}\mathbf{R}^{\prime},$ (22)
-where $\mathbf{R}^{\prime}$ is the rotation matrix derived from non-rigid deformed Gaussian’s rotation quaternion $\mathbf{r}^{\prime}$ and $\mathbf{R_{o}}$ represents the rotation matrix in observation space.
+where $\mathbf{R}^{\prime}$ is the rotation matrix derived from the non-rigid deformed Gaussian’s rotation quaternion $\mathbf{r}^{\prime}$, and $\mathbf{R_{o}}$ represents the rotation matrix in observation space.
 Following the previous method [22], we utilize an MLP $\mathcal{E}_{lbs}$ to predict the LBS weight offsets for each query canonical point and update the sampled weights from the nearest SMPL vertex as
@@
 $I=\mathbf{Splatting}(\mathbf{x_{o}},\mathbf{R_{o}},s^{\prime},\alpha,sh).$ (24)
-During the optimization process, we employ four loss functions as supervision, summarized as
+During the optimization process, we employ a combination of loss functions as supervision, summarized as
 $\mathcal{L}=\lambda_{1}\mathcal{L}_{color}+\lambda_{2}\mathcal{L}_{ssim}+\lambda_{3}\mathcal{L}_{lpips}+\mathcal{L}_{mask},$ (25)
@@
 $\mathcal{L}_{mask}$ [22] is an $L_{2}$ loss between the rendered $\alpha$ and human body mask.
-Additional details of the pipeline are provided in the supplementary material.
+Similar to [56], we apply $\mathcal{L}_{isopos}$ and $\mathcal{L}_{isocov}$ to control the Gaussian primitive’s position and covariance.
+Details are shown in the supplementary material.
 5 Experiments
 5.1 Datasets
-DNA-Rendering [6] is a challenging realistic human-centric rendering dataset.
+DNA-Rendering [9] is a challenging human-centric rendering dataset.
 It contains diverse scenes from everyday life to professional occasions.
-We choose 6 typical sequences with loose-fitting garments (1_0206_04, 2_0007_04, 2_0019_10, 2_0044_11, 2_0051_09, 2_0813_05) for experiments.
+We use 6 sequences with loose-fitting garments (1_0206_04, 2_0007_04, 2_0019_10, 2_0044_11, 2_0051_09, 2_0813_05) for experiments.
 We adopt 24 views for training and 6 views for testing.
-Novel view rendering results compared with other baselines are reported.
-I3D-Human [5] comprises multi-view video frames of humans with loose clothing and complex movements.
-Compared to previous datasets such as ZJU-MoCap [43], which are mainly collected under controlled speeds and tight-fitting garments, I3D-Human contains scenes closer to daily life.
-We conduct experiments on 4 sequences following the data split in Dyco [5] and all training and testing data are retained to be the same across all comparison methods.
-For each sequence, we use 4-5 views for training and the rest for testing.
+I3D-Human [8] contains scenes closer to daily life, which comprises multi-view video frames of humans with loose clothing and complex movements.
+We conduct experiments on 4 sequences and use 4-5 views for training and the rest for testing.
+All training and testing data are retained to be the same in comparisons following Dyco’s [8] data split.
 Table 1: Quantitative Results on DNA-Rendering Dataset.
 We report the performance of novel view rendering.
-Our GAST outperforms previous state-of-the-art human modeling methods.
+Our method outperforms previous state-of-the-art human modeling methods.
 We mark the best and the second best methods in cells.
@@
 […]
-3DGS-Avatar [44] 26.68 0.9434 48.31 28.10 0.9487 53.86 32.09 0.9705 28.87 28.28 0.9506 35.68 25.06 0.9530 41.39 31.75 0.9692 29.94
+3DGS-Avatar [56] 26.64 0.9439 50.09 28.06 0.9492 56.70 32.06 0.9710 31.12 28.23 0.9512 37.15 25.04 0.9534 42.22 31.75 0.9701 31.30
 GART [31] 27.19 0.9454 56.47 28.22 0.9525 56.46 31.13 0.9686 35.16 28.90 0.9561 41.27 26.47 0.9628 42.21 32.06 0.9729 35.74
-GauHuman [17] 28.49 0.9534 46.45 28.32 0.9501 52.16 32.55 0.9716 27.91 31.02 0.9670 28.85 26.24 0.9610 38.21 33.79 0.9776 25.14
-GAST(Ours) 30.81 0.9649 39.47 29.63 0.9566 43.72 34.97 0.9777 23.53 32.62 0.9747 23.83 28.45 0.9685 34.06 35.81 0.9843 20.83
-Figure 3: Qualitative Results on DNA-Rendering Dataset [6].
-We compare the visual quality of novel view rendering with previous SOTA methods.
-To present the differences more clearly, we zoom into the local region and compute the error maps compared with ground truth images.
+GauHuman [22] 27.96 0.9500 50.42 27.93 0.9496 56.93 31.63 0.9698 32.82 30.38 0.9641 34.19 25.99 0.9599 42.00 33.43 0.9764 29.42
+Ours 30.81 0.9649 39.47 29.63 0.9566 43.72 34.97 0.9777 23.53 32.62 0.9747 23.83 28.45 0.9685 34.06 35.81 0.9843 20.83
+Figure 3: Novel View Qualitative Results on DNA-Rendering.
+We zoom into the local region and compute the error maps compared with ground truth images.
 The results show that our method achieves competitive results on both the overall and local region qualities.
-Table 2: Quantitative Results on I3D-Human Dataset.
-We report the performance of novel view rendering.
-The proposed method achieves competitive results compared with Dyco [5], GauHuman [17], and 3DGS-Avatar [44].
+Table 2: Novel View Rendering Quantitative Results on I3D-Human.
 Subject: ID1  1 ID1  2 ID2  1 ID3  1
@@
 3DGS-Avatar [56] 29.78 0.9600 31.40 31.29 0.9627 28.09 29.63 0.9604 37.33 32.73 0.9602 39.47
-Dyco [5] 31.53 0.9648 25.17 31.76 0.9646 25.60 29.95 0.9651 29.46 32.78 0.9631 33.06
+Dyco [8] 30.81 0.9617 28.29 31.45 0.9628 25.75 29.41 0.9618 33.26 32.50 0.9612 35.46
 GauHuman [22] 29.47 0.9571 39.46 30.42 0.9585 38.87 28.51 0.9546 49.46 32.10 0.9546 53.67
-GAST(Ours) 31.81 0.9652 28.64 31.88 0.9636 30.01 31.35 0.9669 34.04 33.47 0.9633 37.90
-Figure 4: Qualitative Results on I3D-Human Dataset [5]. We compare the visual quality of novel view rendering with SOTA methods.
+Ours 31.81 0.9659 27.03 31.98 0.9660 27.84 31.53 0.9685 30.23 33.62 0.9651 34.02
+Figure 4: Novel Pose Qualitative Results on I3D-Human.
+We compare the proposed method with previous SOTA approaches.
+The rendered images and error maps demonstrate our robustness on novel pose rendering.
 5.2 Baselines and Metrics
@@
 Dyco is an implicit NeRF-based method that adopts pose sequence information for temporal modeling.
-We compare performance with Dyco to showcase the better efficacy of our point-based fine-grained spatio-temporal context design.
+We compare performance with Dyco to showcase the better efficacy of our fine-grained motion condition design.
 3DGS-Avatar, GART, and GauHuman are explicit 3DGS-based methods, and all of them are originally designed for monocular inputs.
 For fairness, we extend them to multi-view input under the same settings.
-We compare these methods to demonstrate that finer rendering details can be achieved with our hierarchical sequential conditions for Gaussian points.
 We report three key metrics: peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM) [66], and learned perceptual image patch similarity (LPIPS) [77].
@@
 Figure 5: Qualitative Ablation Results. We compare the visual influence of different components on ID2_1 scene of I3D-Human dataset.
+Figure 6: Out-of-distribution Pose Animation.
 Comparison on DNA-Rendering.
@@
 1 shows the quantitative results on DNA-Rendering dataset.
-The proposed GAST outperforms previous state-of-the-art human modeling methods across all metrics.
+The proposed method outperforms previous state-of-the-art human modeling methods across all metrics.
 To demonstrate the visual improvement, we compare the quality of rendering images in Fig.
@@
 Besides, the results in Fig.
-3 show that the compared methods fail to capture the proper shape variation caused by movement (see the white dress in scene 2_0051_09).
-Although these methods achieve a smooth appearance in these areas, the error maps show that they fail to closely match the actual shape.
-We leverage the spatio-temporal context to predict Gaussian deformations, allowing our method to capture more accurate detailed appearance variations caused by human motions.
+3 show that the compared methods fail to capture the proper shape variation caused by movement (e.g., the white dress in scene 2_0051_09).
+Although these methods achieve a smooth appearance in these areas, the error maps show that they fail to match the actual shape closely.
+We leverage the hierarchical motion context to predict Gaussian deformations, allowing our method to capture detailed appearance variations caused by human motions more accurately.
 Comparison on I3D-Human.
@@
 Previous NeRF-based Dyco [8] utilizes pose variation as a condition to model human motions.
-However, it is limited to capturing only overall body motion trends and lacks the capacity for finer-grained modeling in local regions.
+However, it is limited to capturing only overall body motions and lacks the capacity for finer-grained modeling in local regions.
 For example, in ID1_1 and ID2_1 shown in Fig.
-4, Dyco fails to model the information in regions far from the human body skeleton, leading to rendering results that do not accurately reflect the true motion.
-Our method designs more fine-grained conditions for the explicit point-based Gaussian representation to excavate both spatial and temporal motion information hierarchically.
-The rendering details in Fig.
-4 demonstrate the improvement benefiting from such design in these local regions.
-More results are shown in supplementary materials.
-Table 3: Quantitative Ablation Results on I3D-Human Dataset. The Vanilla Non-Rigid MLP in (b) only utilizes Gaussian positions $\mathbf{x}$ and the current frame pose $P$ as the conditions.
+8 (in the Supp.
+Mat.), our method renders results that accurately reflect the true motion, while Dyco fails to model the information in regions far from the human body skeleton.
+Tab.
+3 reports the quantitative results of novel pose rendering on the I3D-Human dataset.
+Our method achieves best performance among 3DGS-based human modeling methods and also enables real-time rendering ($\sim$ 45 FPS on I3D-Human) compared to NeRF-based Dyco ($\sim$ 0.7 FPS).
+Fig.
+4 shows the qualitative comparisons of novel pose rendering.
+Different from Dyco [8], which relies solely on joint motions as conditions, our method leverages the motion state of each Gaussian primitive to predict deformations, enabling more flexible modeling of motions in regions distant from the human body.
+To evaluate the generalizability, we also conduct experiments on ZJU-MoCap [53]. Please refer to the supplementary materials for more details.
+Out-of-Distribution Poses.
+To evaluate the generalization ability to novel poses, we conduct experiments using animations with large pose variations.
+Specifically, we use a model trained on one sequence to render target poses sampled from other unseen sequences, as different sequences in DNA-Rendering and I3D-Human contain distinct motion patterns.
+As illustrated in Fig.
+6, the Target Pose is drawn from an unseen sequence exhibiting different human motions.
+The results suggest that our method is capable of generalizing to these significant pose changes, even though such poses were not observed during training.
+Table 3: Metrics of Novel Pose Renderings on I3D-Human.
 Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS* $\downarrow$
-(a) Baseline 29.88 0.9570 38.12
-(b) (a)+Vanilla Non-Rigid MLP. 31.18 0.9614 35.78
-(c) (b)+$\Delta\mathcal{P}$ Conditions. 31.89 0.9640 33.86
-(d) (c)+ $\mathcal{V}$ Conditions. 32.05 0.9646 32.82
-(e) (d)+Multi-scale Sampling. (Ours) 32.13 0.9647 32.65
+Dyco [8] 30.09 0.9569 35.32
+3DGS-Avatar [56] 29.54 0.9555 38.59
+GauHuman [22] 29.31 0.9526 48.29
+Ours 30.28 0.9583 36.07
 5.4 Ablation Study
@@
 5 (d) show that our method achieves higher rendering details in local regions.
-Multi-scale sequence sampling is designed to capture both the overall motion trend and the fine-grained inter-frame motion details.
-With temporal conditions sampled at different intervals, our method is able to predict each Gaussian’s deformation more robustly, leading to better rendering quality as shown in Tab.
+Spatio-temporal Multi-scale Sampling (STMS) is designed to enhance the robustness of human motion conditions by incorporating local region motion information and motion patterns across different temporal windows.
+With such a hierarchical motion context embedding design, our method is able to predict each Gaussian’s deformation more robustly, leading to better rendering quality as shown in Tab.
 4 (e) and Fig.
 5 (e).
+Please refer to the supplementary materials for more detailed ablation experiments.
+Table 4: Quantitative Ablation Results on I3D-Human.
+(a) denotes experiments without non-rigid deformation.
+(b) only utilizes Gaussian positions $\mathbf{x}$ and the current frame pose $P$ as the conditions for non-rigid MLP.
+Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS* $\downarrow$
+(a) Baseline 29.76 0.9569 38.35
+[…]
+[…]
+(d) (c)+ $\mathcal{V}$ Conditions. 32.01 0.9651 31.23
+(e) (d)+STMS. (Ours) 32.24 0.9664 29.78
 6 Limitations and Conclusion
 Limitations.
-Compared to prior NeRF-based approaches, our method models the human body with the point-based Gaussian representation, which may introduce needle-like artifacts during rendering.
-As indicated in Tab.
-2, the metric on LPIPS is slightly inferior to that of NeRF-based Dyco [5].
-We notice that some subsequent works [19, 15] on 3DGS optimize the artifacts by introducing surface-related smooth regularization.
-Generative model-based human avatar [26] could also potentially be introduced as a guarantee of high-fidelity rendering.
-However, these are beyond the scope of our motion modeling and will serve as future work to improve rendering quality.
+The Gaussian-based representation used in our method may introduce slight blur artifacts during rendering, whereas NeRF’s ray-based volumetric integration tends to produce sharper results.
+Moreover, our local velocity cues are derived from the coarse SMPL model rather than dense surface tracking [79], which may limit the accuracy of fine-scale garment deformation.
+Addressing these limitations remains an open challenge for future work.
 Conclusion.
-In this paper, we propose GAST, a 3DGS-based framework that integrates hierarchical spatial-temporal information for 3D human modeling.
-Specifically, we introduce a coarse-to-fine motion context for the non-rigid deformation of Gaussian primitives.
-At the spatial dimension, the global and local region transformations for Gaussians are learned based on the overall inter-frame human pose variations and point-wise motions.
-To capture stable temporal motion information for each point, we introduce an SMPL vertex velocity template and sample the velocity of each Gaussian point from it.
-Along the temporal dimension, we apply a multi-scale sampling strategy to capture both long-term motion trends and inter-frame details, supporting more robust predictions of Gaussian deformations.
-Through the above design, we have achieved state-of-the-art rendering of human avatars with complex garments and motions.
+In this paper, we propose a 3DGS-based framework that integrates hierarchical motion context for 3D human modeling.
+Specifically, the non-rigid deformation for the Gaussian primitive is learned based on the global human skeleton variations and fine-grained Gaussian’s point-wise motions.
+To capture each Gaussian primitive’s motion information more robustly, we introduce a spatio-temporal multi-scale sampling strategy to incorporate both local region motion features and motion patterns across different temporal intervals.
+Through the above design, the proposed method achieves state-of-the-art rendering of human avatars with complex garments and motions.
+Acknowledgment
+This research was supported by Shanghai Artificial Intelligence Laboratory.
 Supplementary Material
 7 Details of Skeleton Motion $\Delta P$
-The pose $P$ is represented as the rotations of all joints in axis-angle form:
+The pose $P$ is represented as the rotation relative to the parent node of all joints:
 $P=\{\theta_{1},\theta_{2},...,\theta_{K}\},$ (26)
@@
 8 Datasets
-I3D-Human Dataset [5] We use images with resolution 512 $\times$ 512 for experiments following Dyco [5].
-DNA-Rendering Dataset [6] We use the original images’ resolution 512 $\times$ 612 for experiments.
-For the dataset split, we sample 24 evenly distributed views from camera IDs 0–48 for training and 6 additional uniformly distributed views from IDs 48–60 for novel view testing.
+I3D-Human Dataset [8]. We use images at a resolution of 512 $\times$512 for experiments and follow Dyco[8] for both training and testing splits.
+DNA-Rendering Dataset [9].
+We use images with a resolution of 512 $\times$ 612 for training and 750 $\times$ 1024 for testing in our experiments.
+For the temporal split, we select 100 consecutive time steps (pose IDs 0–99) and use the multi-view images of all selected steps for training.
+During testing, we evaluate on a total of 20 time steps sampled every 5 steps from this training range, using multi-view images from held-out camera views.
+For the view split, we uniformly sample 24 views from camera IDs 0–48 for training and 6 additional views uniformly sampled from camera IDs 48–60 for novel view evaluation.
+ZJU-MoCap [53]. We use images with resolution 512 $\times$ 512 for experiments and follow 3DGS-Avatar[56] for both training and testing splits.
 9 Motion Embeddings for Non-Rigid MLP
-The embeddings $f_{\Delta\mathcal{P}}$ and $f_{\mathcal{V}}$ in Eq.
-(12) describe the skeleton and local region motions, respectively.
-For each frame, all Gaussian primitives $\{\mathcal{G}_{i}\}$ share the same skeleton motion embedding $f_{\Delta\mathcal{P}}\in\mathbb{R}^{32}$ and we concat it with $\mathcal{G}_{i}$’s point-wise motion embedding ${f_{\mathcal{V}}}_{i}\in\mathbb{R}^{64}$ as a new condition $f_{i}^{\prime}\in\mathbb{R}^{96}$.
-$f_{i}^{\prime}$ is then input to a 2-layer MLP $\mathcal{E}_{non-rigid}$ to predict the corresponding non-rigid deformation.
+The embeddings $f_{\Delta\mathcal{P}}$ and $f_{\mathcal{V}}$ in Eq. (17) describe the skeleton and local region motions, respectively.
+For each frame, all Gaussian primitives $\{\mathcal{G}_{i}\}$ share the same skeleton motion embedding $f_{\Delta\mathcal{P}}\in\mathbb{R}^{32}$ and we concat it with $\mathcal{G}_{i}$’s point-wise motion embedding ${f_{\mathcal{V}}}_{i}\in\mathbb{R}^{96}$ as a new condition $f_{i}^{\prime}\in\mathbb{R}^{128}$.
+$f_{i}^{\prime}$ is then input to an MLP $\mathcal{E}_{non-rigid}$ to predict the corresponding non-rigid deformation.
+$\mathcal{E}_{non-rigid}$ consists of 3 hidden layers and one output layer, and each hidden layer is followed by a ReLU activation.
 10 Details of Loss Functions
@@
 $\mathcal{L}_{color}=|I-I_{gt}|.$ (30)
-We use $\mathcal{L}_{ssim}$ to constrain the structure similarity between rendered image and the ground truth, which is given by
+We use $\mathcal{L}_{ssim}$ to constrain the structure similarity between the rendered image and the ground truth, which is given by
 $\mathcal{L}_{ssim}=1-\text{SSIM}(I,I_{gt}),$ (31)
-where $\text{SSIM}(\cdot)$ is the SSIM metric. Additionally, we use the LPIPS loss to ensure the perceptual similarity:
+[…]
 $\mathcal{L}_{lpips}=\text{LPIPS}(I,I_{gt}),$ (32)
-where $\text{LPIPS}(\cdot)$ is the LPIPS metric.
-Following GauHuman [17] and 3DGS-Avatar [44], we employ a mask loss to ensure that Gaussian primitives are accurately localized within their designated regions:
+[…]
 $\mathcal{L}_{mask}={||M-M_{gt}||}_{2}$ (33)
@@
 $M=\sum\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).$ (34)
-Figure 6: Additional Novel View Qualitative Results on DNA-Rendering Dataset [6].
-We provide both the complete image and localized error map comparisons against other methods for a more comprehensive visualization.
+Similar to [56, 54], we also utilize the as-isometric-as-possible constraint [28] to enforce neighboring distance similarity of 3D Gaussian centers and covariance matrices between canonical space and observation space:
+$\mathcal{L}_{isopos}$ $=\sum_{i=1}^{N}\sum_{j\in\mathbf{KNN}(i)}|d(\mathbf{x}^{i},\mathbf{x}^{j})-d(\mathbf{x_{o}}^{i},\mathbf{x_{o}}^{j})|,$ (35)
+$\mathcal{L}_{isocov}$ $=\sum_{i=1}^{N}\sum_{j\in\mathbf{KNN}(i)}|d(\mathbf{\Sigma}^{i},\mathbf{\Sigma}^{j})-d(\mathbf{\Sigma_{o}}^{i},\mathbf{\Sigma_{o}}^{j})|,$ (36)
+where $N$ is the number of Gaussian primitives.
+$\mathbf{x},\mathbf{\Sigma}$ and $\mathbf{x_{o}},\mathbf{\Sigma_{o}}$ are the Gaussian primitive’s position and covariance matrix in canonical space and observation space, respectively.
+$\mathbf{KNN}$ denotes the $K=5$ nearest points.
+Figure 7: Novel View Qualitative Results on DNA-Rendering. We provide both the complete image and localized error map comparisons against other methods for a more comprehensive visualization.
 11 Details of Optimization
 In our experiments, the number $\lvert\mathcal{S}\rvert$ (in Eq.
-(16)) of sequence sampled in different scales is set to 3 and each sequence length $L$ (in Eq.
+(14)) of sequence sampled in different scales is set to 3, and each sequence length $L$ (in Eq.
 (6)) is set to 8.
@@
 Additionally, we utilize the adaptive densification in  [27] to control the number of Gaussian primitives.
-For DNA-Rendering dataset, we use the SMPL-X [42] model with pose dimension $P\in\mathbb{R}^{25\times 3}$ in our experiments.
+For the DNA-Rendering dataset, we use the SMPL-X [51] model with pose dimension $P\in\mathbb{R}^{25\times 3}$ in our experiments.
 We adopt the SMPL-X template with $N=10475$ vertexes as Gaussian initialization.
@@
 (14)).
-For I3D-Human dataset, we follow Dyco [5] to use SMPL [37] model in our experiments for fairness.
+For the I3D-Human dataset, we follow Dyco [8] to use the SMPL [43] model in our experiments for fairness.
 The number of SMPL template vertexes used for Gaussian initialization is $N=6890$ and the corresponding pose dimension is $P\in\mathbb{R}^{23\times 3}$.
@@
 4 (d) adopts single sampling step that $\mathcal{S}=\{s_{0}\}$.
+Figure 8: Novel View Qualitative Results on I3D-Human.
 12 Additional Results
-12.1 Comparisons of Novel View Rendering
-Figure 7: Additional Novel Pose Qualitative Results on I3D-Human Dataset [5].
+12.1 Experiments on ZJU-MoCap
+ZJU-MoCap [53] is a common benchmark dataset in human avatar, which is mainly collected under controlled speeds and tight-fitting garments.
+We use 6 sequences (377, 386, 387, 392, 393, 394) for experiments following the dataset split in 3DGS-Avaar[56].
+Comparison on ZJU-MoCap.
+Results on DNA-Rendering and I3D-Human demonstrate that our methods achieve better performance in complex cases with loose clothing and uncontrolled speeds.
+Furthermore, Tab.
+5 and Fig.
+9 show that our method also achieves competitive metrics in simpler controlled scenarios, which verifies the generalizability of the proposed methods.
+Table 5: Quantitative Results on ZJU-MoCap.
+Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS* $\downarrow$
+Dyco [8] 30.37 0.9599 29.47
+3DGS-Avatar [56] 30.59 0.9609 27.09
+GauHuman [22] 31.04 0.9620 31.81
+Ours 31.02 0.9619 28.89
+Figure 9: Qualitative Results on ZJU-MoCap.
+12.2 Visual Comparisons On DNA-Rendering
 Fig.
@@
 Our approach excels in regions where motion causes notable appearance variations, producing results that conform to the true shape more accurately, as exemplified by the sleeve (top row) and the skirt hem (bottom row) in Fig.
-6.
-12.2 Comparisons of Novel Pose Rendering
-Table 4: Quantitative Results of Novel Pose Renderings on I3D-Human Dataset.
-Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS* $\downarrow$
-3DGS-Avatar [44] 29.53 0.9554 38.58
-Dyco [5] 30.30 0.9582 33.52
-GauHuman [17] 29.31 0.9525 48.29
-GAST(Ours) 30.39 0.9582 38.15
-Tab.
-4 reports the quantitative results of novel pose rendering on the I3D-Human dataset [5].
-Our method achieves competitive performance compared with SOTA human modeling methods.
-Fig.
-7 shows the qualitative comparisons of novel pose rendering and we utilize error maps to visualize the differences between various methods more clearly.
-Compared to Dyco [5], which relies solely on joint motions as conditions, our GAST leverages the motion state of each Gaussian primitive to predict deformations, enabling more flexible modeling of motions in regions distant from the human body.
-The shape of the clothing in Fig.
-7 demonstrates that our method can render the appearance consistent with the novel pose motion, leading to more accurate results.
+7.
+12.3 Video Quality
+We provide rendered videos on the DNA-Rendering dataset [9], along with the corresponding error maps similar to Fig.
+3.
+Please refer to the attached videos for comparison.
+In the videos, the novel view results are rendered on temporal frames with pose IDs 0–99, using camera viewpoints that are held out during training.
 13 Additional Ablations
@@
 Sampling Step $\mathcal{S}$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS* $\downarrow$
-$\mathcal{S}=\{24\}$ 32.05 0.96468 32.82
-$\mathcal{S}=\{33\}$ 32.04 0.96467 33.06
-$\mathcal{S}=\{42\}$ 32.06 0.96472 32.87
-$\mathcal{S}=\{24,33,42\}$ 32.13 0.96476 32.65
+$\mathcal{S}=\{24\}$ 32.01 0.9655 30.71
+$\mathcal{S}=\{33\}$ 32.09 0.9656 30.88
+$\mathcal{S}=\{42\}$ 32.11 0.9658 30.81
+$\mathcal{S}=\{24,33,42\}$ 32.24 0.9664 29.78
 Different KNN Sampled Vertexes Tab.
 7 presents the quantitative results on I3D-Human dataset [8] with different numbers of KNN vertexes when deriving the velocity $e_{i}$ of each Gaussian primitive (Eq.
-11).
+12).
 It suggests that considering the motion states of all local regions helps obtain a more robust Gaussian primitive’s velocity embedding, contributing to enhancing performance.
-Table 6: Impact of Different KNN Vertex Number $\tau$.
-KNN num. $\tau$ $\tau=1$ $\tau=3$ $\tau=6$ $\tau=8$
-PSNR $\uparrow$ 32.06 32.07 32.08 32.13
+Table 7: Impact of Different KNN Vertexe Number $\tau$.
+KNN num. $\tau$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS* $\downarrow$
+$\tau=1$ 32.14 0.9659 30.41
+$\tau=5$ 32.15 0.9660 30.32
+$\tau=8$ 32.24 0.9664 29.79
+$\tau=10$ 32.20 0.9663 30.11
+14 Computational Cost Analysis
+Table 8: Computational Efficiency.
+Mehods Train Time FPS Train Mem.
+Dyco 6 h 0.7 20 GB
+3DGS-Avatar 25 m 25 6 GB
+Ours ($3k$ $iter$) 5 m 62 8 GB
+We provide computational cost analysis for experiments on ZJU-Mocap (conducted on a single 4090 GPU).
+Tab.
+8 shows that our approach achieves competitive computational efficiency.
+The additional computation time for each frame’s Vertex Motion Template (multi-scale sampling number $|\mathcal{S}|=3$, sequence length $L=8$) is about 0.09 s, and it only needs to be computed once during preprocessing before training.